# Tadashi: Enabling AI-Based Automated Code Generation With Guaranteed Correctness

### v1

@@ Abstract.
Frameworks and DSLs auto-generating code have traditionally relied on human experts developing them to have in place rigorous methods to assure the legality of the applied code transformations. Machine Learning (ML) is gaining wider adoption as a means to auto-generate code optimised for the hardware target. However, ML solutions, and in particular black-box DNNs, provide no such guarantees on legality. In this paper we propose a library, Tadashi, which leverages the polyhedral model to empower researchers seeking to curate datasets crucial for applying ML in code-generation. Tadashi provides the ability to reliably and practically check the legality of candidate transformations on polyhedral schedules applied on a baseline reference code. We provide a proof that our library guarantees the legality of generated transformations, and demonstrate its lightweight practical cost. Tadashi is available at https://github.com/vatai/tadashi/.
@@ 1. Introduction
Machine learning (ML) is disrupting all forms of digital media, and the creation of computer programs is no exception. ML can be and has been applied in various ways with the aim of making the work of programmers (Dehaerne et al., 2022a) easier: generating code from “natural languages” (Phan, 2019), write a summary (e.g., write commit messages) of a piece of code (Shido et al., 2019), auto-complete code on-the-fly while it’s being written (Github, [n. d.]) etc. In this paper, we consider the task of optimising code to a hardware target, i.e., using ML to transform the hotspots in the code, i.e., loops, in a way that improves performance.
Checking if the program generated by the ML model is faster then the baseline is just the matter of running it, then comparing against the run-time of the original code. However, it is a major challenge to get formal guarantees that the transformed code produces the same output as the original (Dehaerne et al., 2022b; Munoz et al., 2024; Jiang et al., 2024). This impedes the adoption of ML for code generation and transformation in mission critical scientific/engineering codes, or codes that come with high liability, such as in the financial, legal, and corporate sectors. Additionally, there is a lack of tooling for generating and sharing parameterised datasets comprised of legal auto-generated code, which further constraints progress in the field since, to date, ML-based solutions have to curate their datasets.
To this end, we developed Tadashi, a Python library for sampling the search space of possible loops transformations, with the capability to verify the legality (i.e., correctness) of the sampled transformations. Tadashi allows the user to programmatically specify the pattern for sampling legal individual or combinations of transformations, based on the preference of the ML method that would be used to explore the search space of transformations. Through the use of the polyhedral model, Tadashi is capable of enabling ML engineers to explore novel and more aggressive methods of code optimisation in the datasets they curate, while guaranteeing that the results of the generated code are functionally identical to the original program. Tadashi key point is the ease-of-use for ML practitioners; Tadashi aims to provide a simple Python interface, which can be used to curate datasets or perform direct code-generation to be used with different ML approaches.
@@
We introduce Tadashi, a Python library designed for ML practitioners seeking to design tools for code auto generation with correctness guarantees. Tadashi lowers the barrier to entry of polyhedral loop transformations; a model that can give mathematically provable guarantees on the correctness (i.e., legality) of the transformed code.
We demonstrate how Tadashi can be used by an ML system by implementing an end-to-end stub reinforcement learning system.
We evaluate performance to confirm it is not an obstacle in Tadashi’s adoption. We observe throughput measurements above 32 transformations and legality checks per second, going up to 781 for some benchmarks. The overheads of parsing and generating code is negligible in comparison to compiling and running the generated code.
@@ 2.1. Correctness in ML-Based Code Auto-Generation
For guarantees on correctness, existing approaches for using ML in code generation or transformation rely on one of the following approaches:
@@
In summary, the approaches above either provide no formal guarantee on functional correctness, or confine the user to a very constrained form of code generation. In addition, we observed that for the above approaches, it was the burden of the researchers to implement a sampling strategy and curate the dataset. Solving those two problems, correctness and agency, is the motivation for Tadashi.
@@ 2.2. The Polyhedral Model
then the schedule is legal, because it maps the dependent to a later time then the timestep of the dependee. Because (2) is equivalent to , we can (symbolically) compute and if is lexicographically positive, the schedule is legal. Computing is useful because it provides us with even more useful information: if an entry in is zero, it means there are no dependencies along the corresponding dimension of the schedule, and consequently the generated loop for that dimension can be parallelised.
@@ 2.3. Constraints of the Polyhedral Model
To be able to represent the domain, schedule, and dependencies efficiently, the source codes represented by the polyhedral model are restricted to Static Control Parts, or SCoPs for short (Listing 1 is a simple example). While these restrictions are not negligible, most of high performance and science codes do not violate them. The main restrictions of SCoPs are the following: (1) The control-flow must be static, which means it does not depend on the input, other than parameters (such as loop boundaries) which remain constant during execution, (2) there is no memory aliasing, no pointer manipulation, and (3) the symbolic expressions and formulas, used throughout the model, need to be expressed using a matrix and a vector as a (quasi-)affine expression, i.e., (sometimes referred to as Presburger formulas (Verdoolaege et al., [n. d.])). For example, the domains are expressed as ; schedules are (quasi-)affine functions . The polyhedral model can also handle tiling, so the formulas include integer division with a constant as well, and this is what the “quasi-” prefix refers to. However, for brevity and clarity, we omit discussion of these details and for a more comprehensive description of the polyhedral model we recommend (Feautrier, 1996, 1997; Verdoolaege et al., [n. d.]).
@@ 3. Tadashi: A Library to Sample Legal Polyhedral Transformations
Tadashi is a library that provide the means for ML practitioners to sample the space of legal polyhedral loop transformations. The user can use Tadashi to describe a sampling strategy that is suited for the ML approach, e.g. uniform sampling to curate datasets vs. guided exploration of the space of legal transformations. Figure 1 gives an overview of Tadashi. At an architectural level, Tadashi consists of a Python front-end and a C/C++ back-end (aptly named ctadashi). The front-end provides a minimal and easy-to-use interface to sample from the space of possible transformations, check their legality, and generate the transformed code. More importantly, the front-end being in Python encourages fast adoption by ML practitioners since Python is the de facto language used in AI/ML frameworks. The back-end implements this functionality, using the Integer Set Library (ISL) (Verdoolaege, 2010) and Polyhedral Extract Tool (PET) (Verdoolaege and Grosser, 2012) polyhedral libraries.
@@ 3.1. Python Front-End: Polyhedral Made Easy
To demonstrate the ease of use of Tadashi, Listing 2 performs a tiling transformation, checks legality, and generates new code in less than 10 lines of a fully functional Python example. Line 2 creates the app object, which stores the information required by Tadashi to load the input file. The scops object in line 3 stores the polyhedral representations of the loop nests. The node object on line 4 specifies the loop inside the first SCoP, and a tiling transformation (with tile size 16) is performed on line 5. The legality of the tiling is checked together with the transformation, and the result is stored in the legal variable. If the transformation is legal, the new code is generated in line 7.
@@
To pass an input file to Tadashi, it must be stored in an instance of the App abstract base class. Currently Tadashi provides two implementations of App: the Simple class (used in line 2) which can be used for simple single-file programs and Polybench (used later in Section 5) to load Polybench apps (Reisinger, 2024). The classes derived from App store the input filename to Tadashi and additionally include directories are also needed by Tadashi when parsing the file. For convenience, App also defines compile() and measure() methods.
@@
The polyhedral representation of all loop nests (SCoPs) are implemented as an array-like scops object (line 3), and from the first SCoP (scops[0]) a node of the schedule trees (Verdoolaege et al., 2014) is selected (line 4). For a user not familiar with the polyhedral model and schedule trees, a node is “an entity that can be transformed”, which usually means a loop in the loop nest (a discussion of schedule trees can be found in Section 4). Since some transformations are not available for certain loops, the Node class provides valid() and valid_args() methods to check availability. The nodes of a scop can be accessed through the (dynamic) scop.schedule_tree[] as a list, or as a tree using the goto_child(child_idx) and goto_parent() methods of nodes.333The root is scop.schedule_tree[0]!.
A code transformation is performed by passing a TrEnum (an Enum type enumerating all transformations) object to node.transform() (with the transformation-specific argument). A list of transformations with a short description of their arguments can be found in Table 1.
These are primitive transformations, in the sense that they are composable, so we can apply a sequence of transformations on the same SCoP. In the current version of Tadashi, we prioritised the the implementation of loop transformations that have high impact on performance (Pouchet et al., 2011)
@@
The return value of the transform() method is the result of the legality check and, based on it, the users can decide what they want to do. In case of a legal transformation, one likely action would be generating the C code (line 7), compiling it and measuring the execution time, while for an illegal transformation one could invoke the rollback() method of the node, to return the polyhedral representation to legal state.
@@
The Python front-end, aims to hide the details of the polyhedral model. To this end, Tadashi provides aliases to class and variable names which might be daunting or confusing to a user unfamiliar with polyhedral concepts (see Table 2). This way, Tadashi can hide the details of what is happening under the hood, which we describe in Section 4.
@@ 3.1.1. Sampling Strategies
Listing 2 demonstrates the application of a single primitive transformation. However, in consequence of the compositionality of polyhedral transformations, the user can apply a sequence of transformations and legality checks to a SCoP (before generating the code) and hence obtain a different composite transformation. This is a new sample from the space of polyhedral transformations, different from its component primitive transformations. Put differently, sampling the space of polyhedral transformations with Tadashi consists of applying a sequence of transformations to a SCoP, where each entry consists of specifying a location (a node in the schedule tree), the transformation (TrEnum) and a set of arguments. A location can be selected using the schedule_tree[] member or with a combination of goto_* methods of the Node class. The transformation can be selected from the possible values TrEnum values and checked if it is available on the selected node. Finally, selecting an argument is somewhat convoluted, since the available values depend not only on the selected transformation, but also on the concrete properties of the selected schedule node (such as its partial schedule, type and number of children etc.). Tadashi exposes the necessary information for random argument selection by providing the list of available values or the lower and upper bounds for integer arguments.
For example, if we intend to narrow the search space by iterating over combinations of shifts and tiles, shifting by single increments of surrounding iteration variables and tile sizes of powers of starting from to , the following steps need to be taken. The desired loop is saved as a node, and each transformation is combined with its arguments. For tiling, this can easily be accomplished using list comprehension, by gathering (TrEnum.TILE, 2**i) pairs where i iterates over range(3, 11). To create the argument list of the shift transformation, the user obtains the lower and upper bounds of all arguments for FULL_SHIFT_VAR on node. The first argument, the coefficient, is set to since we want single increments but the second argument, the iterator index is varied, i.e., list of triplets is created containing FULL_SHIFT_VAR, 1 and iter_idx, where iter_idx iterater over the range between the lower and upper bound of the second argument. Using the product() function from Python’s itertools package, these two lists are combined into a list of (tile, shift) pairs. Then by iterating through these pairs in a loop, we can call the node’s transform() method, first with *tile, then with *shift, followed by a checking the legality and generating code.
| Polyhedral | Simplified |
Scops |
LoopNests |
Scop |
LoopNest |
scop.schedule_tree[] |
loop_nest.nodes[] |
@@ 3.2.2. Supervised Learning
Such approach is especially valuable, because evaluation of transformed code takes significantly more time, compared to selecting and performing transformations (see sec. 5)
The choice of transformations can be done in random or by grid search. Many transformations require additional corresponding parameters, e.g. choosing to tile a loop requires choosing a tile size. These arguments (second column in Table 1) allow the user to combinatorially increase the coverage of the search space to consequently increase the coverage of the curated dataset. For example, starting with a limited set of input C source files containing loop nests, one could run Tadashi to select random transformations (on a random input file, on random nodes of the schedule tree, with random parameters), then store the new code with the result of the legality check and the measured execution time. The dataset generated this way would consist of (input, output) pairs, where the input is (the name of) a C source file, and the output consists of: (1) the transformation, (2) transformed code, (3) legality of the transformation, (4) execution time of the transformed program.
@@
3.2.3. Evolutionary Algorithms
Similar to RL, evolutionary algorithms use a mix of exploration and exploitation to traverse the search space, however without relying on on learnable parameters and performing search using more direct heuristics (Cai et al., 2023). Exploration refers to searching for new, diverse solutions by exploring a wide range of the solution space, while exploitation focuses on refining and improving existing solutions by intensively searching around the best solutions found so far. As mentioned above in other ML approaches, Tadashi’s sampling of the search space can be parameterised by the user to explore a wide range of the search space (exploration) or intensive search in a narrow region of the search space (exploitation). The runtime of the transformed code variants would be used as the objective function.
The exploration capability of evolutionary algorithms can in some instances improve if infeasible solutions appear in the population (While and Hingston, 2013), and then later be penalised to disappear over time from the population (Coello, 2022). Tadashi internally manipulates the schedule to generate both legal and illegal transformations, where we then filter out the illegal transformations. Since illegal transformations can be of value for the exploration capability of evolutionary algorithms, we allow the user to sample illegal transformations simply by not rolling them back and generating code disregarding the result of the legality check.

### v2

@@ Abstract.
Frameworks and domain-specific languages for auto-generating code have traditionally depended on human experts to implement rigorous methods ensuring the legality of code transformations. Recently, machine learning (ML) has gained traction for generating code optimized for specific hardware targets. However, ML approaches-particularly black-box neural networks-offer no guarantees on the correctness or legality of the transformations they produce. To address this gap, we introduce Tadashi, an end-to-end system that leverages the polyhedral model to support researchers in curating datasets critical for ML-based code generation. Tadashi provides an end-to-end system capable of applying, verifying, and evaluating candidate transformations on polyhedral schedules with both reliability and practicality. We formally prove that Tadashi guarantees the legality of generated transformations, demonstrate its low runtime overhead, and showcase its broad applicability. Tadashi available at https://github.com/vatai/tadashi/.
@@ 1. Introduction
Machine learning (ML) is disrupting all forms of digital media, and the creation of computer programs is no exception. ML can be and has been applied in various ways with the aim of making the work of programmers (Dehaerne et al., 2022) easier: generating code from “natural languages” (Phan, 2019), write a summary (e.g., write commit messages) of a piece of code (Shido et al., 2019), auto-complete code on-the-fly while it’s being written (Github, 2022) among others. In this paper, we consider the task of optimising code to a hardware target, i.e., using ML to transform the hotspots in the code, i.e., loops, in a way that improves performance.
Checking if the program generated by the ML model is faster then the baseline is just a matter of running it, then comparing against the run-time of the original code. However, it is a major challenge to get formal guarantees that the transformed code produces the same output as the original (Dehaerne et al., 2022; Munoz et al., 2024; Jiang et al., 2024). This impedes the adoption of ML for code generation and transformation in mission critical scientific/engineering codes, or codes that come with high liability, such as in the financial, legal, and corporate sectors. Additionally, there is a lack of tooling for generating and sharing parameterized datasets comprised of legal auto-generated code, which further constraints progress in the field since, to date, ML-based solutions have to curate their datasets.
To this end, we developed Tadashi, a Python library for sampling the search space of possible loop transformations, with the capability to verify the legality (i.e., correctness) of the sampled transformations. Tadashi allows the user to programmatically specify the pattern for sampling legal individual or combinations of transformations, based on the preference of the ML method that would be used to explore the search space of transformations. Through the use of the polyhedral model, Tadashi is capable of enabling ML engineers to explore novel and more aggressive methods of code optimisation in the datasets they curate, while guaranteeing that the results of the generated code are functionally identical to the original program. Tadashi is an end-to-end solution to transform, verify and evaluate transformations, emphasizing the ease-of-use for ML practitioners, with support for HPC oriented libraries such as MPI4Py futures (Rogowski et al., 2023); Tadashi aims to provide a simple Python interface, curating datasets or performing direct code-generation at scale to be used with different ML approaches.
We envision Tadashi with the overarching purpose of helping clusters and supercomputers reach their full potential. Namely, after the deployment of a new (or existing) supercomputing system, one can train an ML model to learn to transform hotspots in scientific software to run optimally on the given hardware. By basing their ML models on Tadashi, end users will have both fully correct and highly optimized code. With this in mind Tadashi would be a viable and practicable option when porting existing science codes to new supercomputers. This is particularly a huge challenge, as demonstrate in recent history by the Exascale Computing Project (Heroux, 2024; exa, [n. d.]), which required 7-years, $1.8 billion, and approximately 1,000 researchers from DOE national laboratories to port the codes to new systems (among other goals of ECP).
@@
We introduce Tadashi, a Python-based end-to-end system designed for ML practitioners seeking to design tools for code auto generation with correctness guarantees. Tadashi lowers the barrier to entry of polyhedral loop transformations; a model that can give mathematically provable guarantees on the correctness (i.e., legality) of the transformed code.
We demonstrate how Tadashi can be used by an ML system by implementing two representative ML models on top of Tadashi.
We show the capability of Tadashi to speed up software on Polybench suite and two science applications.
@@ 2.1. Correctness in ML-Based Code Auto-Generation
To ensure correctness, existing approaches for using ML in code generation or transformation rely on one of the following approaches:
@@
In summary, the approaches above either provide no formal guarantee on functional correctness, or confine the user to a very constrained form of code generation. While some have suggested there is a tradeoff between correctness and performance, (Ouyang et al., 2024) we take a different view—emphasizing that correctness remains paramount. This perspective is reinforced by recent findings (202, 2025) where what was initially reported as a 100x speedup from an AI-generated solution was, upon closer examination, actually a 3x slowdown. In addition, we observed that for the above approaches, it was the burden of the researchers to implement a sampling strategy and curate the dataset. Solving those two problems, correctness and agency, is the motivation for Tadashi.
@@ 2.2. The Polyhedral Model
then the schedule is legal, because it maps the dependent to a later time then the timestep of the dependee. Because (2) is equivalent to , we can (symbolically) compute and if is lexicographically positive, the schedule is legal. Computing is useful because it provides us with even more useful information: if an entry in is zero, it means there are no dependencies along the corresponding dimension of the schedule, and consequently the generated loop for that dimension can be parallelized.
@@ 2.3. Constraints of the Polyhedral Model
To be able to represent the domain, schedule, and dependencies efficiently, the source codes represented by the polyhedral model are restricted to Static Control Parts, or SCoPs for short (Listing 1 is a simple example). While these restrictions are not negligible, most of high performance and science codes do not violate them. The main restrictions of SCoPs are the following: (1) The control-flow must be static, which means it does not depend on the input, other than parameters (such as loop boundaries) which remain constant during execution, (2) there is no memory aliasing, no pointer manipulation, and (3) the symbolic expressions and formulas, used throughout the model, need to be expressed using a matrix and a vector as a (quasi-)affine expression, i.e., (sometimes referred to as Presburger formulas (Verdoolaege et al., 2021)). For example, the domains are expressed as ; schedules are (quasi-)affine functions . The polyhedral model can also handle tiling, so the formulas include integer division with a constant as well, and this is what the “quasi-” prefix refers to. However, for brevity and clarity, we omit discussion of these details and for a more comprehensive description of the polyhedral model we recommend (Feautrier, 1996, 1997; Verdoolaege et al., 2021).
@@ 3. Tadashi: A Library to Sample Legal Polyhedral Transformations
Tadashi is a library that provide the means for ML practitioners to sample the space of legal polyhedral loop transformations. The user can use Tadashi to describe a sampling strategy that is suited for the ML approach, e.g., uniform sampling to curate datasets vs. guided exploration of the space of legal transformations. Figure 1 gives an overview of Tadashi. At an architectural level, Tadashi consists of a Python front-end and a C/C++ back-end (aptly named ctadashi). The front-end provides a minimal and easy-to-use interface to sample from the space of possible transformations, check their legality, and generate the transformed code. More importantly, the front-end being in Python encourages fast adoption by ML practitioners since Python is the de facto language used in AI/ML frameworks. The back-end implements this functionality, using the Integer Set Library (ISL) (Verdoolaege, 2010) and Polyhedral Extract Tool (PET) (Verdoolaege and Grosser, 2012) polyhedral libraries.
@@ 3.1. Python Front-End: Polyhedral Made Easy
To demonstrate the ease of use of Tadashi, Listing 2 performs a loop fission, checks legality, generates new code and measures speedup in no more than 10 lines of a fully functional Python example. Line 3 creates the app object, which stores the information required by Tadashi to load the input file. The location of the transformation is determined by selecting a node objects on line 4. The fission transformation is performed on line 5. The new code code is generated in line 6.
@@
To pass an input file to Tadashi, it must be wrapped in an instance of the App abstract base class. Currently Tadashi provides two implementations of App: the Simple class which can be used for simple single-file programs and Polybench (used in line 3) to load Polybench kernels (Reisinger, 2024). The classes derived from App store the input filename to Tadashi and other information needed by Tadashi when parsing the file. For convenience, App also defines compile() and measure() methods.
@@
The polyhedral representation of all loop nests (SCoPs) are implemented as an array-like scops object (line 4), and from the first SCoP (scops[0]) a node of the schedule trees (Verdoolaege et al., 2014) is selected. For a user not familiar with the polyhedral model and schedule trees, a node is “an entity that can be transformed”, which usually means a loop in the loop nest (a brief discussion of schedule trees can be found in Section 4). Since some transformations are not available for certain loops (e.g., it doesn’t make sense to apply 3D tiling on a 2-deep loop nest), the Node class provides methods to query the available transformations and their (valid) arguments. The nodes of a scop can be accessed through the (dynamic) scop.schedule_tree[] as a list, or as a tree using the goto_child(child_idx) and goto_parent() methods of nodes.333The root is scop.schedule_tree[0]!.
A code transformation is performed by passing a TrEnum (an Enum type enumerating all transformations) object to the node’s transform() method (with the transformation-specific argument). A summary of implemented transformations (and their parameters) can be found in Table 1. TILE1D, (TILE2D and TILE3D) implements 1D, (2D and 3D) tiling (higher dimentional tiling can be expressed tiling and INTERCHANGE transfromations). INTERCHANGE switches an outer loop with the loop directly under it. For PARTIAL shift transformations, the statement index (stmt idx) selects only one statement to be shifted, while the FULL versions shift all statements. The VAL shift transformation shift the schedule by a constant, the VAR and PARAM versions shift by an iterator (iter idx) of an outer loop or a parameter (param idx), multiplied by a coefficient (coeff). FULL_FUSE fuses all loops in a sequence, while FUSE fuses two selected loops. Similarly, FULL_SPLIT and SPLIT perform loop fission. SET_PARALLEL and SET_LOOP_OPT modify how code of a loop is generated. These are primitive transformations, in the sense that they are composable, so we can apply a sequence of transformations on the same SCoP. In the current version of Tadashi, we prioritized the the implementation of loop transformations that have high impact on performance (Pouchet et al., 2011).
@@
The return value of the transform() method is the result of the legality check and, based on it, the users can decide what they want to do. In case of a legal transformation, one likely action would be generating the C code (line 6 in Listing 2), compiling it (line 8) and measuring the execution time (line 9), while for an illegal transformation one could invoke the rollback() method of the node, to return the polyhedral representation to the previous, legal state. Additionally, Scop objects provide a reset method and a method to apply a list of transformations, useful to implement “arbitrary rollback”.
@@
FULL_FUSE |
SPLIT |
1 stamt index |
FULL_SPLIT |
@@ 3.1.1. Sampling Strategies
Listing 2 demonstrates the application of a single primitive transformation. However, in consequence of the compositionality of polyhedral transformations, the user can apply a sequence of transformations and legality checks to a SCoP (before generating the code) and hence obtain a different composite transformation. This is a new sample from the space of polyhedral transformations, different from its component primitive transformations. Put differently, sampling the space of polyhedral transformations with Tadashi consists of applying a sequence of transformations to a SCoP, where each entry consists of specifying a location (a node in the schedule tree), the transformation (TrEnum) and a set of arguments. A location can be selected using the schedule_tree[] member or with a combination of goto_* methods of the Node class. The transformation can be selected from the possible values TrEnum values and checked if it is available on the selected node. Finally, selecting an argument is somewhat convoluted, since the available values depend not only on the selected transformation, but also on the concrete properties of the selected schedule node (such as its partial schedule, type and number of children etc.). The get_args() method of Nodes simplifies this.
@@ 3.2.2. Supervised Learning
Such approach is especially valuable, because evaluation of transformed code takes significantly more time, compared to selecting and performing transformations (see Section 5.4).
The choice of transformations can be done in random or by grid search. Many transformations require additional corresponding parameters, e.g., choosing to tile a loop requires choosing a tile size. These arguments (second column in Table 1) allow the user to combinatorially increase the coverage of the search space to consequently increase the coverage of the curated dataset. For example, starting with a limited set of input C source files containing loop nests, one could run Tadashi to select random transformations (on a random input file, on random nodes of the schedule tree, with random parameters), then store the new code with the result of the legality check and the measured execution time. The dataset generated this way would consist of (input, output) pairs, where the input is (the name of) a C source file, and the output consists of: (1) the transformation, (2) transformed code, (3) legality of the transformation, (4) execution time of the transformed program.
@@
3.2.3. Evolutionary Computation Algorithms
Similar to RL, evolutionary computation (EC) algorithms use a mix of exploration and exploitation to traverse the search space, however without relying on learnable parameters and performing search using more direct heuristics (Cai et al., 2023). Exploration refers to searching for new, diverse solutions by exploring a wide range of the solution space, while exploitation focuses on refining and improving existing solutions by intensively searching around the best solutions found so far. As mentioned above in other ML approaches, Tadashi’s sampling of the search space can be parameterized by the user to explore a wide range of the search space (exploration) or intensive search in a narrow region of the search space (exploitation). The runtime of the transformed code variants would be used as the objective function.
The exploration capability of EC algorithms can in some instances improve if infeasible solutions appear in the population (While and Hingston, 2013), and then later be penalized to disappear over time from the population (Coello, 2022). Tadashi internally manipulates the schedule to generate both legal and illegal transformations, where we then filter out the illegal transformations. Since illegal transformations can be of value for the exploration capability of EC algorithms, we allow the user to sample illegal transformations simply by not rolling them back and generating code disregarding the result of the legality check.